OpenStudy (umulas):

Can anyone explain to me how a partial fraction decomposition work? I got a question of 18 -------- x(x^2+6)

OpenStudy (fibonaccichick666):

So for partial fractions you essentially factor your denominator and create two fractions from what was before, 1. So in general, it is of the form \[\frac{f(x)}{some factor-able function}=\frac{A}{1st Factor}+\frac{B}{2nd Factor}\] Now when you have a factor that is a polynomial, you must include one less than that above. So with your quadratic, the B will become Bx+C.

OpenStudy (fibonaccichick666):

So, if we apply this to your problem. The bottom is already factored for us. So we set it up as such:\[\frac{18}{x(x^2+6)}=\frac{A}{x}+\frac{Bx+C}{x^2+6}\]

OpenStudy (fibonaccichick666):

So now, what do we have to do in order to add two fractions?

OpenStudy (umulas):

multiply both sides, right?

OpenStudy (fibonaccichick666):

Not quite, think back to when you first learned how to add \[\frac{1}{3}+\frac{1}{2}\] What did you have to do?

OpenStudy (umulas):

oh make sure the other side is equal in the denominator 1/3 + 1/2 -> 2/6 + 3/6 = 5/6

OpenStudy (fibonaccichick666):

There we go. So we had to find a common denominator. To do that we had to multiply by something that reduced to 1 right?

OpenStudy (fibonaccichick666):

In the example it was 2/2 or 3/3 yea?

OpenStudy (umulas):

I'm sorry i'm confused; i don't know what to do for the numerator

OpenStudy (fibonaccichick666):

ok, so how did you turn 1/3 into 2/6?

OpenStudy (umulas):

just multiplied by 2

OpenStudy (johnweldon1993):

Wel, you multiplied by \(\large \frac{2}{2}\) right?

OpenStudy (umulas):

right

OpenStudy (fibonaccichick666):

if you only multiplied by 2, you would have 2/3 yea?

OpenStudy (umulas):

yes

OpenStudy (fibonaccichick666):

So, the same principle applies here. Your denominator is x, so what do you have to multiply the other fraction by in order to create a common denominator?

OpenStudy (fibonaccichick666):

\[\frac{18}{x(x^2+6)}=\frac{A}{x}(\frac{?}{?})+\frac{Bx+C}{x^2+6}(\frac{?}{?})\] We know our common denominator is \(x(x^2+6)\). We just need to turn the right into one fraction.

OpenStudy (umulas):

A(x^2+6)+(Bx+C) =========== x(x^2+c)

OpenStudy (fibonaccichick666):

Close, you forgot to multiply the second fraction by something

OpenStudy (fibonaccichick666):

well, reallly the second numerator

OpenStudy (johnweldon1993):

And also watch that 'c' in the denominator should be 6 :)

OpenStudy (fibonaccichick666):

I'm also assuming the c on the bottom is a type-o

OpenStudy (umulas):

A(x^2+6)+(Bx+C)x =========== x(x^2+6) ??

OpenStudy (fibonaccichick666):

There we go.

OpenStudy (fibonaccichick666):

So now get rid of all of your parentheses and group the like terms

OpenStudy (umulas):

including the numerator?

OpenStudy (fibonaccichick666):

Sorry, I should have specified, only the numerator

OpenStudy (fibonaccichick666):

Don't worry about the denominator at the moment

OpenStudy (umulas):

Ax^2+6A + Bx^2 +Cx ------------------- x(x^2+6)

OpenStudy (fibonaccichick666):

Good, so now we have : \[\frac{18}{x(x^2+6)}=\frac{Ax^2+6A+Bx^2+Cx}{x(x^2+6)}\]

OpenStudy (fibonaccichick666):

So, notice how the denominators are the same on both sides?

OpenStudy (fibonaccichick666):

That means that the tops must be the same thing

OpenStudy (fibonaccichick666):

so, we don't have an x^2 term on the left right?

OpenStudy (fibonaccichick666):

or an x for that matter, yea?

OpenStudy (johnweldon1993):

*In terms of the numerator...remember the denominators are now gone \[\large \frac{18}{\cancel{x(x^2+6)}}=\frac{Ax^2+6A+Bx^2+Cx}{\cancel{x(x^2+6)}}\] \[\large 18 = Ax^2 + 6A + Bx^2 + Cx\]

OpenStudy (umulas):

hmm ok but does the letter A nad B stay?

OpenStudy (fibonaccichick666):

So, here is where we use logic

OpenStudy (fibonaccichick666):

since you don't have an x^2 or an x term on the left, what must their coefficients be?

OpenStudy (umulas):

3?

OpenStudy (fibonaccichick666):

If you multiply x^2 by 3 it will go away?

OpenStudy (fibonaccichick666):

What can you multiply something by to make it disappear?

OpenStudy (umulas):

0?

OpenStudy (fibonaccichick666):

yea, so we really have \[0x^2+0x+18=Ax^2+6A+Bx^2+Cx\]

OpenStudy (umulas):

I'm not sure this is going the right direction, since my answer got 3 -3x - + ----- x x^2 + 6

OpenStudy (fibonaccichick666):

That is entirely correct

OpenStudy (umulas):

so how do we reach that?

OpenStudy (fibonaccichick666):

I thought you reached that yourself?

OpenStudy (umulas):

No, in my test, I'm just practicing for a final and wanted to know how i can do that

OpenStudy (fibonaccichick666):

oh, ok, so we are here:\[0x^2+0x+18=Ax^2+6A+Bx^2+Cx\]

OpenStudy (fibonaccichick666):

Now, we can set the like terms equal to each other and solve. So for example, we have: \[0x=Cx\] When we solve for C, we get C=0. Right?

OpenStudy (umulas):

right

OpenStudy (fibonaccichick666):

We can do the same for each type of term. So, you try setting it up for the constant term

OpenStudy (fibonaccichick666):

Don't solve, just set it up first

OpenStudy (umulas):

set what up?

OpenStudy (fibonaccichick666):

Set up the equation for the constant terms

OpenStudy (fibonaccichick666):

Like I did for the x terms

OpenStudy (umulas):

You mean A + Bx+C --- --------- x^2+6

OpenStudy (fibonaccichick666):

Not quite. So above we had: "Now, we can set the like terms equal to each other and solve. So for example, we have: 0x=Cx When we solve for C, we get C=0. Right?"

OpenStudy (fibonaccichick666):

We set the x terms equal and solved for C, I would like you to do this for the constant terms(the ones without an x).

OpenStudy (fibonaccichick666):

So, if you are confused, start by telling me which terms are constant.

OpenStudy (johnweldon1993):

\[0x^2+0x+18=Ax^2+6A+Bx^2+Cx\] *Set the \(\large x^2\) terms on the right side equal to the \(\large x^2\) terms on the left side* etc...

OpenStudy (fibonaccichick666):

You still there? @umulas

OpenStudy (fibonaccichick666):

Just in case I am not on when you come back, you only have like 3 more steps. Here is a good resource for you: http://tutorial.math.lamar.edu/Classes/CalcII/PartialFractions.aspx